# Effective Community Detection Over Streaming Bipartite Networks (Technical Report)

### v1

@@ Abstract.
The streaming bipartite graph is extensively used to model the dynamic relationship between two types of entities in many real-world applications, such as movie recommendations, location-based services, and online shopping. Since it contains abundant information, discovering the dense subgraph with high structural cohesiveness (i.e., community detection) in the bipartite streaming graph is becoming a valuable problem. Inspired by this, in this paper, we study the structure of community on the butterfly motif in the bipartite graph. We propose a novel problem, named Community Detection over Streaming Bipartite Network (CD-SBN), which aims to retrieve qualified communities with user-specific query keywords and high structural cohesiveness at snapshot and continuous scenarios. In particular, we formulate the user relationship score in the weighted bipartite network via the butterfly pattern and define a novel $(k,r,\sigma)$-bitruss as the community structure. To efficiently tackle the CD-SBN problem, we design effective pruning strategies to rule out false alarms of $(k,r,\sigma)$-bitruss and propose a hierarchical synopsis to facilitate the CD-SBN processing. Due to the dynamic of streaming bipartite networks, we devise an efficient procedure for incremental graph maintenance. We develop an efficient algorithm to answer the snapshot and continuous CD-SBN query by traversing the synopsis and applying the pruning strategies. With extensive experiments, we demonstrate the efficiency and effectiveness of our proposed CD-SBN processing approach over real/synthetic streaming bipartite networks.
@@
PVLDB, 14(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@
(a) $G_{t-1}$ at timestamp $(t-1)$
(b) $G_{t}$ at timestamp $t$
Figure 1. An example of community detection in streaming click bipartite networks $G_{t}$.
@@ 1. Introduction
Nowadays, the management of bipartite graphs that involve two distinct types of nodes (e.g., customers and products, visitors and check-in locations, users and clicked websites) has been extensively studied in many real applications such as Online Recommendation from user-product transaction graphs (Fan et al., 2022; Weng et al., 2022; Kou et al., 2023), location-based services with visitor-location check-in graphs (Hang et al., 2018; Zhou et al., 2021a; Wang et al., 2017; Kim et al., 2022), and behavior analysis in user-Webpage click graphs (Sun et al., 2009; Neria et al., 2017).
The community detection (CD) over such bipartite graphs has recently become increasingly important for bipartite graph analysis and mining (Barber, 2007; Bouguessa and Nouri, 2021; Yen and Larremore, 2020; Zhou and Amini, 2019), which uncovers hidden and structurally cohesive bipartite subgraphs (or groups). The CD problem over bipartite graphs has many useful real applications, such as online product recommendation and marketing, user behavior analysis, malicious user group identification, and so on.
Although many prior works (Wang et al., 2023; Abidi et al., 2023a; Wang et al., 2021; Zhang et al., 2024) studied the CD problem in static bipartite graphs, real-world bipartite graphs are usually dynamically changing over time (e.g., updates of purchase transactions, check-ins, or website browsing). Therefore, in this paper, we will consider the Community Detection over Streaming Bipartite Network (CD-SBN), upon graph updates (e.g., edge insertions/deletions or edge weight changes).
Below, we give a motivating example of our CD-SBN problem in the application of detecting malicious user groups (communities) over a user-website click bipartite graph.
@@
(Anomaly Detection for the Cybersecurity) Figure 1(a) illustrates an example of a clickstream bipartite network $G_{t-1}$ at timestamp $(t-1)$, which contains 3 user vertices, $u_{1}\sim u_{3}$, 2 website vertices, $v_{1}$ and $v_{2}$, and 5 edges $(u_{i},v_{a})$ between two types of nodes, user $u_{i}$ and website $v_{a}$ (for $1\leq i\leq 3$ and $1\leq a\leq 2$). Here, each website vertex $v_{a}$ (for $a=1,2$) has a set of keywords that represent the website’s features (e.g., {bank, finance} for a bank website $v_{1}$). Moreover, each edge $(u_{i},v_{a})$ (for $i=1\sim 3$ and $a=1,2$) is associated with an integer weight, indicating the frequency that user $u_{i}$ accessed website $v_{a}$ for the past five minutes.
@@
As shown in Figure 1(a), to detect malicious groups of users accessing sensitive/suspicious websites, the network security officer may want to specify some query keywords of websites (e.g., “bank” and “malware”), and identify some group (community), $g_{1}$, of users (e.g., $u_{2}$ and $u_{3}$) who frequently visit some common websites with query keywords (e.g., bank website $v_{1}$ and hacking website $v_{2}$). The officer will warn those suspicious users in the resulting community $g_{1}$ and/or take immediate actions (e.g., recording the evidence for the police).
Figure 1(b) shows the streaming bipartite network $G_{t}$ at timestamp $t$, with a newly inserted edge $(u_{1},v_{1})$ and an expired edge $(u_{3},v_{2})$ (i.e., user $u_{3}$ has not visited website $v_{2}$ for the past five minutes). In this case, at timestamp $t$, the community in $G_{t}$ is changed to $g_{2}$ (rather than $g_{1}$ at timestamp $(t-1)$), as circled by the orange dashed line. $\blacksquare$
@@
In addition to the example above, the CD-SBN problem has many other real applications. For example, in the customer-product transaction bipartite graph, we can use the CD-SBN results to find communities of customers who recently have a purchasing behaviors for online advertising and marketing. Similarly, in the visitor-location check-in bipartite graph (e.g., from Yelp (Inc., 2024)), we can identify a group of visitors who frequently check in at some common points of interest (POIs) and provide them with group buying coupons/discounts (e.g., Groupon (Groupon, 2024)).
@@
In this paper, we formulate and tackle the CD-SBN problem, which obtains all bipartite communities with the user-specified query keywords (e.g., POIs’ features) and high structural cohesiveness (e.g., small distance between any two users and high structural score of the community). In particular, we formally define the community semantics in the context of bipartite graphs, that is, a keyword-aware and structurally dense bipartite subgraph (i.e., a so-called $(k,r,\sigma)$-bitruss that contains query keywords, as will be described in Section 2.4). We consider both snapshot and continuous scenarios of our CD-SBN problem, which detect communities over a snapshot of streaming bipartite network $G_{t}$ at timestamp $t$, or continuously monitor the CD-SBN answer set for each registered community constraint upon streaming graph updates, respectively.
@@
Due to the large scale of bipartite graphs and rapid streaming graph changes, efficient processing of the snapshot and continuous CD-SBN is rather challenging. In order to tackle the challenges, in this paper, we propose effective pruning strategies with respect to the community constraints (e.g., query keywords, community radius, edge supports, and community scores) to significantly reduce the CD-SBN problem space. We design a hierarchical synopsis to effectively facilitate the candidate community search, and develop efficient snapshot and continuous algorithms to retrieve or incrementally maintain actual community answers (via our proposed synopsis and pruning methods), respectively.
@@
    We present a framework to efficiently process the two types of queries with incremental updates of the bipartite streaming graph in Section 3.
@@
    We develop an efficient algorithm with effective synopsis-level pruning strategies to answer the snapshot CD-SBN query and another algorithm to maintain the result set of a continuous CD-SBN query with low computational cost in Section 6.
@@ 2.1. Streaming Bipartite Networks
For simplicity, this paper considers vertices $u$ in the upper vertex layer $U(G)$ as users in real-world applications, and vertices $v$ in the lower vertex layer $L(G)$ as items that are associated with keyword sets $v.K$ such as movie types, POI features, and product descriptions. Moreover, each edge $e_{u,v}$ from user $u$ to item $v$ is associated with a weight $w_{u,v}$, indicating the user-item interaction frequency such as ##\## of movie views for movie recommendations, ##\## of POI visits by users, ##\## of product purchases for online shopping, and so on.
@@
(Streaming Bipartite Network, $G_{t}$) A streaming bipartite network $G_{t}$ consists of an initial bipartite graph $G_{0}=(U(G_{0}),L(G_{0}),E(G_{0}),\Phi(G_{0}))$, and an ordered sequence of update items $S=\{p_{1},p_{2},\ldots,p_{t},\ldots\}$, where each item $p_{t}=(e_{u,v},t)$ indicates an update operator of weight $w_{u,v}$ for an inserted edge $e_{u,v}$ arriving at timestamp $t$. A sliding window, $W_{t}$, of size $s$ in a streaming bipartite network $G_{t}$ contains the most recent $s$ update items $p_{t-s+1}$, $p_{t-s+2}$, $\ldots$, and $p_{t}$, over bipartite graph $G_{t}$, where $t$ is the current timestamp.
In Definition 2.2, we consider a sliding window, $W_{t}$, of $s$ update items over a streaming bipartite network $G_{t}$. Each update item can be the insertion of a new edge/interaction between users and items (or equivalently the increase of the edge weight $w_{u,v}$).
After applying update items in sliding window $W_{t}$ to the bipartite graph $G$, we can obtain the latest snapshot, $G_{t}$, of the streaming bipartite network at timestamp $t$.
@@
Insertion of a New Item $p_{t+1}$. For the insertion of a new item $p_{t+1}$ $=(e_{u,v},t+1)$, we consider two cases:
@@ 2.2. Basic Patterns in the Bipartite Graph
Figure 2. An example of basic patterns over bipartite network.
@@
Based on Definition 2.3, a wedge $\angle(u_{i},v,u_{j})$ in the bipartite graph can be two audiences watching the same movie, two travelers visiting the same site, or two customers purchasing the same product.
The weight of the wedge $\angle(u_{i},v,u_{j})$ takes the minimum edge weights between $w_{u_{i},v}$ and $w_{u_{j},v}$, indicating the intensity (or frequency) of maximally common interest/behavior $v$ (e.g., movie, POI site, or product) between two users $u_{i}$ and $u_{j}$.
In the example of Figure 2, $\angle(u_{1},v_{1},u_{2})$ is one of the wedges, with weight $w_{\angle(u_{1},v_{1},u_{2})}=\min\{2,4\}=2$.
@@
The butterfly structure (as given in Definition 2.4) has been widely used in many algorithms and applications (Sanei-Mehri et al., 2018; Xu et al., 2022b; Zhou et al., 2021b) to detect communities with high cohesiveness in bipartite graphs. In Figure 2, $\Join(u_{2},u_{3},v_{2},v_{3})$ is an example of a butterfly.
@@
$\vspace{-3ex}w_{\Join(u_{i},u_{j},v_{a},v_{b})}$ $=$ $w_{\angle(u_{i},v_{a},u_{j})}\cdot w_{\angle(u_{i},v_{b},u_{j})}$
@@ 2.3. The $(k,r,\sigma)$-Bitruss Community in a Bipartite Graph
In this subsection, we first define the user relationship score between any two user vertices $u_{1}$ and $u_{2}$ in a bipartite graph $G$ and then formalize the $(k,r,\sigma)$-bitruss community with high structural cohesiveness in $G$.
The Relationship Score of a User Vertex Pair: Based on the score of a butterfly (as given in Eq. (2.5)), we can define the user relationship score of a user vertex pair $(u_{1},u_{2})$ as follows.
@@
[…]
@@
where $dist(u_{c},u_{i})$ is the shortest path distance between $u_{c}$ and $u_{i}$ in bipartite subgraph $g$.

### v2

@@ Abstract.
The streaming bipartite graph is widely used to model the dynamic relationship between two types of entities in various real-world applications, including movie recommendations, location-based services, and online shopping. Since it contains abundant information, discovering the dense subgraph with high structural cohesiveness (i.e., community detection) in the bipartite streaming graph is becoming a valuable problem. Inspired by this, in this paper, we study the structure of the community on the butterfly motif in the bipartite graph. We propose a novel problem, named Community Detection over Streaming Bipartite Network (CD-SBN), which aims to retrieve qualified communities with user-specific query keywords and high structural cohesiveness at snapshot and continuous scenarios. In particular, we formulate the user relationship score in the weighted bipartite network via the butterfly pattern and define a novel $(k,r,\sigma)$-bitruss as the community structure. To efficiently tackle the CD-SBN problem, we design effective pruning strategies to rule out false alarms of $(k,r,\sigma)$-bitruss and propose a hierarchical synopsis to facilitate the CD-SBN processing. We develop efficient algorithms to answer snapshot and continuous CD-SBN queries by traversing the synopsis and applying pruning strategies. With extensive experiments, we demonstrate the performance of our CD-SBN approach on real/synthetic streaming bipartite networks.
@@
PVLDB, 19(1): XXX-XXX, 2026.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
@@
(a) $G_{t-1}$ at time (t1)(t-1)
(b) $G_{t}$ at time $t$
(c) $G_{t+1}$ at time $(t+1)$
Figure 1. A community detection example in streaming click bipartite networks $G_{t}$.:
@@ 1. Introduction
Nowadays, the management of bipartite graphs that involve two distinct types of nodes (e.g., customers and products, visitors and check-in locations, users and clicked websites) has been extensively studied in many real applications such as online recommendation from user-product transaction graphs (Fan et al., 2022; Weng et al., 2022; Kou et al., 2023), location-based services with visitor-location check-in graphs (Hang et al., 2018; Zhou et al., 2021; Wang et al., 2017; Kim et al., 2022), and behavior analysis in user-webpage click graphs (Sun et al., 2009; Neria et al., 2017).
The community detection (CD) over such bipartite graphs has recently become increasingly important for bipartite graph analysis and mining (Barber, 2007; Bouguessa and Nouri, 2021; Yen and Larremore, 2020; Zhou and Amini, 2019), which uncovers hidden and structurally cohesive bipartite subgraphs (or groups). The CD problem over bipartite graphs has numerous useful real-world applications, including online product recommendations and marketing, user behavior analysis, and identifying malicious user groups.
Although many prior works (Wang et al., 2023; Abidi et al., 2023; Wang et al., 2021; Zhang et al., 2024; Wang et al., 2022; Zhang et al., 2021) studied the CD problem in static bipartite graphs, real-world bipartite graphs are usually dynamically changing over time (e.g., updates of purchase transactions, check-ins, or website browsing). Therefore, in this paper, we will consider the Community Detection over Streaming Bipartite Network (CD-SBN), upon graph updates (e.g., edge insertions/deletions or edge weight changes). Below, we provide a motivating example of our CD-SBN problem to detect malicious user groups (communities) on a user-website click bipartite graph.
@@
(Anomaly Detection for the Cybersecurity) In the application of online anomaly detection, the network security officer may want to specify query keywords for websites (e.g., “bank” and “malware”) and identify a group (community) of malicious users accessing sensitive websites. Users in communities frequently visit common websites with query keywords (e.g., bank or hacking websites). The officer will warn suspicious users in the resulting community and take immediate actions (e.g., record evidence for the police).
Figure 1(a) illustrates an example of a clickstream bipartite network $G_{t-1}$ at timestamp (t1)(t-1), which contains 3 user vertices, $u_{1}\sim u_{3}$, 2 website vertices, $v_{1}$ and $v_{2}$, and 5 edges $(u_{i},v_{a})$ between user $u_{i}$ and website $v_{a}$ (for $1\leq i\leq 3$ and $1\leq a\leq 2$). Here, each website vertex $v_{a}$ has a set of keywords that represent the website’s features (e.g., {bank, finance} for a bank website $v_{1}$). Moreover, each edge $(u_{i},v_{a})$ is associated with an integer weight, indicating the frequency that user $u_{i}$ accessed website $v_{a}$ over the past five minutes. At timestamp (t1)(t-1), the officer notices a community $g_{1}$ (circled by blue dashed lines) and starts to monitor suspicious users $u_{2}$ and $u_{3}$ in $g_{1}$.
At timestamp $t$, a new access by user $u_{1}$ to website $v_{1}$ (i.e., edge $(u_{1},v_{1})$) is recorded. Meanwhile, since the officer only focuses on suspicious users who frequently visit sensitive websites within five minutes (rather than normal users who visit infrequently over a long period), access records five minutes ago will expire (i.e., edge $(u_{3},v_{2})$ is deleted). Due to this edge deletion, community $g_{1}$ is no longer cohesive (i.e., less suspicious), so the officer will pay less attention to $g_{1}$.
Finally, at timestamp $(t+1)$, due to frequent accesses by user $u_{1}$ to website $v_{1}$ and the expiration of the access by user $u_{3}$ to $v_{1}$, the officer will pinpoint a new community $g_{2}$ (as circled by orange dashed lines in Figure 1(c)) and take actions against users and websites in $g_{2}$.  $\blacksquare$
@@
In addition to the example above, the CD-SBN problem has many other real applications. For example, in the customer-product transaction bipartite graph, we can utilize the CD-SBN results to identify communities of customers who have recently exhibited purchasing behaviors related to online advertising and marketing. Similarly, in the visitor-location check-in bipartite graph (e.g., from Yelp (Inc., 2024)), we can identify a group of visitors who frequently check in at some common points of interest (POIs) and provide them with group buying coupons/discounts (e.g., Groupon (Groupon, 2024)).
@@
In this paper, we formally define the community semantics in the context of dynamic bipartite graphs, that is, a keyword-aware and structurally dense bipartite subgraph (called $(k,r,\sigma)$-bitruss containing query keywords, as described in Section 2.4). Then, we will tackle the CD-SBN problem, which identifies all bipartite communities with user-specified vertex features and high structural cohesiveness. We consider both snapshot and continuous scenarios for our CD-SBN problem. The former detects communities over a snapshot of the streaming bipartite network $G_{t}$ at timestamp $t$, whereas the latter continuously monitors CD-SBN answers for each registered community constraint upon streaming graph updates.
@@
Due to the large scale of bipartite graphs and the rapid streaming of graph changes, efficiently processing both the snapshot and the continuous CD-SBN is rather challenging. Thus, we propose effective pruning strategies regarding community constraints (e.g., query keywords, community radius, edge supports, and community scores) to significantly reduce the CD-SBN problem space. We also design a hierarchical synopsis to effectively facilitate candidate community search and develop efficient snapshot and continuous algorithms to retrieve or incrementally maintain actual community answers (via proposed synopsis and pruning methods), respectively.
@@
    We present an efficient CD-SBN framework for processing queries over streaming bipartite graphs in Section 3.
@@
    We develop an efficient algorithm with effective synopsis-level pruning strategies to answer the snapshot CD-SBN query, as well as another algorithm to maintain the result set of a continuous CD-SBN query with low computational cost, in Section 6.
@@ 2.1. Streaming Bipartite Networks
For simplicity, this paper considers vertices $u$ in the upper vertex layer $U(G)$ as users in real-world applications, and vertices $v$ in the lower vertex layer $L(G)$ as items that are associated with keyword sets $v.K$ such as movie types, POI features, and product descriptions. Moreover, each edge $e_{u,v}$ from user $u$ to item $v$ is associated with a weight $w_{u,v}$, indicating the user-item interaction frequency such as #\# of movie views for movie recommendations, #\# of POI visits by users, #\# of product purchases for online shopping, and so on.
@@
(Streaming Bipartite Network, $G_{t}$) A streaming bipartite network $G_{t}$ consists of an initial bipartite graph $G_{0}=(U(G_{0}),L(G_{0}),E(G_{0}),\Phi(G_{0}))$, and an ordered sequence of update items $S=\{p_{1},p_{2},\ldots,p_{t},\ldots\}$, where each item $p_{t}=(e_{u,v},t)$ indicates an update operator of weight $w_{u,v}$ for an inserted edge $e_{u,v}$ arriving at timestamp $t$. A sliding window, $W_{t}$, of size $s$ in a streaming bipartite network $G_{t}$ contains the most recent $s$ update items $p_{t-s+1}$, $p_{t-s+2}$, $\ldots$, and $p_{t}$, in $G_{t}$, where $t$ is the current timestamp.
In Definition 2.2, we consider a sliding window, $W_{t}$, of $s$ update items over a streaming bipartite network $G_{t}$. In reality, streaming updates are interactions between users and items in real-world applications (e.g., users’ ratings of movies, access records of POIs, and purchase records). Each update can be the insertion of a new edge/interaction between a user and an item (or equivalently, the increase of the edge weight $w_{u,v}$). Since the interaction records are time-sensitive, we employ a sliding window, $W_{t}$, to focus on recently coming edges and ignore those expired edge updates.
After applying update items in the sliding window $W_{t}$ to the bipartite graph $G$, we can obtain the latest snapshot, $G_{t}$, of the streaming bipartite network at timestamp $t$.
@@
Insertion of a New Item $p_{t+1}$. For the insertion of a new item $p_{t+1}$ $(e_{u,v},t+1)$, we consider two cases:
@@ 2.2. Basic Patterns in the Bipartite Graph
Based on Definition 2.3, a wedge $\angle(u_{i},v,u_{j})$ in the bipartite graph can be two audiences watching the same movie, two travelers visiting the same site, or two customers purchasing the same product. The weight of the wedge $\angle(u_{i},v,u_{j})$ takes the minimum edge weights between $w_{u_{i},v}$ and $w_{u_{j},v}$, indicating the intensity (or frequency) of maximally common interest/behavior $v$ (e.g., movie, POI site, or product) between two users $u_{i}$ and $u_{j}$. In the example of Figure 2, $\angle(u_{1},v_{1},u_{2})$ is one of the wedges, with weight $w_{\angle(u_{1},v_{1},u_{2})}=\min\{2,4\}=2$.
Figure 2. An example of basic patterns over a bipartite network.:
@@
The butterfly structure (as given in Definition 2.4) has been widely used in many algorithms and applications (Sanei-Mehri et al., 2018; Xu et al., 2022b; Zhou et al., 2021) to detect communities with high cohesiveness in bipartite graphs. In Figure 2, $\Join(u_{2},u_{3},v_{2},v_{3})$ is an example of a butterfly.
@@
$\vskip-12.91663ptw_{\Join(u_{i},u_{j},v_{a},v_{b})}$ $=$ $w_{\angle(u_{i},v_{a},u_{j})}\cdot w_{\angle(u_{i},v_{b},u_{j})}$
@@ 2.3. The $(k,r,\sigma)$-Bitruss Community in a Bipartite Graph
In this subsection, we first define the user relationship score between any two user vertices $u_{i}$ and $u_{j}$ in a bipartite graph $G$ and then formalize the $(k,r,\sigma)$-bitruss community with high structural cohesiveness in $G$.
The Relationship Score of a User Vertex Pair: Based on the score of a butterfly (as given in Eq. (2.5)), we can define the user relationship score of a user vertex pair $(u_{i},u_{j},u_{j})$ as follows.
@@
As illustrated in Figure 2, the relationship score, $score_{u_{2},u_{3}}$, between users $u_{2}$ and $u_{3}$ is given by: $w_{\Join(u_{2},u_{3},v_{1},v_{2})}+w_{\Join(u_{2},u_{3},v_{1},v_{3})}+w_{\Join(u_{2},u_{3},v_{2},v_{3})}=\min\{4,2\}\cdot\min\{5,1\}+\min\{4,2\}\cdot\min\{6,3\}+\min\{5,1\}\cdot\min\{6,3\}=2\times 1+2\times 3+1\times 3=11$.
@@
where $u_{c}\in g$, and $dist(u_{c},u_{i})$ is the shortest path distance between $u_{c}$ and $u_{i}$ in bipartite subgraph $g$.
The $(k,r,\sigma)$-bitruss $g$ in Definition 2.7 is defined as a connected subgraph that: 1) each edge is contained in at least $k$ butterflies, 2) subgraph $g$ is centered at $u_{c}$ with a maximum radius $2r$, and 3) each pair of user vertices $u_{i},u_{j}$ share a common neighbor $v$ holding $score_{u_{i},u_{j}}(g)\geq\sigma$. Intuitively, in real applications, the $(k,r,\sigma)$-bitruss refers to a user group with strong connections (measured by a user relationship score), for example, users who enjoy movies of the same genre, visitors who visited the same POIs before, or customers with many common purchased products.